OpenStudy (anonymous):

3y=-1/2x+2 y=-x+9 use substitution

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3(-x+9)=-1/2x+2 -3x+27=-1/2x+2 +(-1/2x) -3/1/2x+27=2 -27 -27 -3/1/2x=-25 x=7/1/7 i believe